# Transcendence: Generative Models Can Outperform The Experts That Train Them

### v3

@@ 3.2 Transcendence with Low-Temperature Sampling
The above shows that, even though transcendence cannot be achieved when directly modeling the distribution, it can be achieved by temperature sampling, assuming that the arg-max predictor achieves higher reward compared to all experts. In other words, we make the subtle connection here that low-temperature sampling can be thought of as performning majority vote [1, 6] between the experts. When the experts put non-negligible mass onto the best actions, the resulting majority vote may find the best action [9], which improves performance compared to individual experts (i.e., “wisdom of the crowd”) and thus achieve transcendence.
@@ 3.3 Denoising a Single Expert
We now turn to study particular cases where low-temperature sampling can lead to transcendence. The most simple case is of a single expert that outputs correct but noisy predictions. Denote by $f^{*}$ the optimal expert, s.t. for all $x$ we have  $f^{*}(y|x)=\dfrac{\delta(y\in Y^{*}_{x})}{\lvert Y^{*}_{x}\rvert}$, where $Y^{*}_{x}=\{y\in\mathcal{Y}:y=\max_{y^{\prime}}r(x,y^{\prime})\}$ and $\delta(\text{condition})$ is 1 if the condition is true and 0 otherwise. Now, for some $\rho\in(0,1)$, let $f_{\rho}$ be a “noisy” expert, s.t., for all $x$, with probability $\rho$ chooses a random output, and with probability $1-\rho$ chooses an output according to the optimal expert $f^{*}(\cdot|x)$, namely $f_{\rho}(y|x)=\rho/\left\lvert\mathcal{Y}\right\rvert+(1-\rho)f^{*}(y|x)$. We show that transcendence is achieved with low-temperature sampling for data generated by $f_{\rho}$:
Theorem 3.
@@ 3.4 Transcendence from Multiple Experts
Next, we consider the case where the dataset is generated by multiple experts that complement each other in terms of their ability to correctly predict the best output. For example, consider the case where the input space is partitioned into $k$ disjoint subsets, $\mathcal{X}=\mathcal{X}_{1}\dot{\cup}\dots\dot{\cup}\mathcal{X}_{k}$, s.t. the $i$-th expert performs well on the subset $\mathcal{X}_{i}$, but behaves randomly on other subsets. Namely, assume the expert $f_{i}$ behaves as follows:  $f_{i}(y|x)=\left(\frac{\delta(x\in\mathcal{X}_{i},y\in Y_{x}^{*})}{\lvert Y^{*}_{x}\rvert}+\frac{\delta(x\notin\mathcal{X}_{i})}{\lvert\mathcal{Y}\rvert}\right)$ where $Y_{x}^{*}$ is as previously defined and $\delta(\text{condition})$ is 1 if the condition is true and 0 otherwise. We show that, assuming that the test distribution $p_{\mathrm{test}}$ is not concentrated on a single subset $\mathcal{X}_{i}$, we achieve transcendence with low-temperature sampling:
Theorem 4.
@@
In order to build intuition for Theorem 4, see Appendix C for an intuitive diagram of the theorem.
@@ 4 Experiments
To evaluate the predictive power of our impossibility result of transcendence with no temperature sampling (Theorem 1) as well as our result of transcendence from multiple experts with low temperature sampling (Theorem 2), we turn to modeling and training chess players. Chess stands out as an attractive option for several reasons. Chess is a well-understood domain and more constrained than other settings such as natural language generation, lending to easier and stronger analysis. Evaluation of skill in chess is also natural and well-studied, with several rigorous statistical rating systems available. In this paper, we use the Glicko-2 rating system [7], which is also adopted by https://lichess.org, the free and open-source online chess server from which we source our dataset.
@@ 4.2 Experimental Results
In this section we attempt to answer our primary research question, can low-temperature sampling actually induce transcendence in practice? We test Theorem 2 by evaluating several ChessFormers across different temperature values, from $0.001$ (nearly deterministic), to $1.0$ (original distribution), to $1.5$ (high entropy). In Figure 1 we definitively confirm the existence of transcendence. Our ChessFormer 1000 (where the latter number refers to the maximum rating seen during training) and ChessFormer 1300 models are able to transcend to around 1500 rating at temperature $\tau$ equal to $0.001$. Interestingly, ChessFormer 1500 is unable to transcend at test time, a result that we further analyze in subsection 4.2.
@@
$F(f^{\prime},f;x)=\mathbb{E}_{x\sim d^{f^{\prime}},y\sim f^{\prime}(\cdot|x)}[r(x,y)]-\mathbb{E}_{x\sim d^{f^{\prime}},y\sim f^{\prime}(\cdot|x)}[r(x,y)].$ (4)
@@
In Figure 4, we find that lowering the temperature has the effect of skewing the expected reward distribution to the left, especially for the green $\tau=0.001$ distribution. This result implies that the model does not improve the expected reward by a small amount for many game states, but rather improves the expected reward by a relatively large amount for a few game states. Thus, $\tau=0.001$ improves the expected reward (probability of winning) by an average of $\mathbf{2.15\pm 0.17\%}$, but for some states, this expected improvement is over 5%. Note that the original temperature expected reward can be thought of as a Dirac distribution centered at 00.  The above finding answers our research question in this section: Low-temperature sampling is able improves the expected reward by relatively large amounts for some specific game states, which is likely why the ChessFormer $1000$ and $1300$ model was able to achieve transcendence.
@@
Importantly, we cannot calculate this normalized entropy for every state, as most states after move $16$ in the midgame and before the engame are unique within the dataset and we therefore observe just a single action for thus states. Therefore our metric is limited in that it only considers opening moves, the beginning of the midgame, and the endgame. We consider only common states with greater than $100$ actions by sampling $1,000,000$ games from each dataset. The average entropy confirm our hypothesis: The $<1500$ cut off dataset has on average less diversity than the $<1300$ dataset, which has is again less than the $<1000$ dataset. This result suggests that Chessformer $1500$ likely is not transcendent due to a lack of diversity in its dataset. If the entropy instead stayed constant for each dataset, it would imply that each had a similar level of diversity. In such a case, we would expect that ChessFormer $1500$ likely would also transcend. Instead, as predicted, Chessformer 1500 likely is not transcendent due to a lack of diversity in its dataset.
@@ 5 Related Work
Another historical thread in AI research is the strength of diverse learners. Long since the development of ensemble methods that exploit learner diversity—including bagging [1], boosting [6], and model averaging [19]—researchers have continued to articulate this insight across settings. Similar to our chess setting, a diverse team of go playing agents have been proven and empirically shown to outperform solitary agents [9] and homogeneous teams [28], even when the alternative models individually outperform the diverse team members [17]. We draw a connection to this deep literature through our theoretical results which shows that training on just the imitation learning objective and then performing low-temperature sampling subtly implies the same principle of majority voting used in this literature.
Teacher diversity has also been explored in the machine learning literature. One related method is ensemble distillation [16], in which a model is trained with an additional objective to match a variety of weaker teacher models. Closer to our setting, ensemble self-training approaches [24] train a learner directly on the labels produced by varied teachers. Large language models supervised by smaller or less trained models are said to exhibit “weak to strong generalization” [2]. Overall, evidence continues to accrue that the general phenomenon we address is pervasive: that is, models can substantially improve over the experts that generate their training data.
@@
Our work also draws connections to the Offline Reinforcement Learning [14] setting, where one attempts to learn a new policy $\pi$ that improves upon a fixed dataset generated by some behavior policy $\pi_{\beta}$. However, our setting of imitation learning differs substantially from this literature, as we do not explicitly train our model on a RL objective that attempts to improve upon the dataset. Importantly, such an objective oftentimes introduces training instabilities [15] and also assumes reward labels, both of which are avoided with a pure imitation or self-supervised learning objective. We defer a more extended discussion of related work to Appendix D.
@@ 6 Discussion and Future Work
This paper introduces the concept of transcendence, where generative models trained on expert data outperform the best individual experts. Our theoretical analysis shows that low-temperature sampling is key to achieving transcendence by denoising expert biases and consolidating diverse knowledge. We validate our findings empirically by training several chess models which, under low-temperature sampling, surpass the performance of the players who produced their training data. We highlight the necessity of dataset diversity for transcendence, emphasizing the role of varied expert perspectives.
@@
The possibility of “superintelligent” AGI has recently fueled many speculative hopes and fears. It is therefore possible that our work will be cited by concerned communities as evidence of a threat, but we would highlight that the denoising effect addressed in this paper does not offer any evidence for a model being able to produce novel solutions that a human expert would be incapable of devising. In particular, we do not present evidence that low temperature sampling leads to novel abstract reasoning, but rather denoising of errors.
@@ Appendix A Proofs
Here we prove Theorem 1, where transcendencecannot occur by purely using imitation learning in our setting where all experts are sampled uniformly across the input distribution.
@@
We now give the proof of Theorem 2 that if the arg-max prediction is better than the best expert, then transcendence is possible with low-temperature sampling.
@@
To prove Theorem 3, we directly use the result in Theorem 2.
@@
Finally, we give the proof Theorem 4, or the statement that transcendence can occur from multiple experts if the test distribution $p_{\mathrm{test}}$ is spread across multiple disjoing subsets of $\mathcal{X}_{i}$.
@@
Thus, we get $f_{\max}=f^{*}$, and the required follows from Theorem 2.

### v4

@@ 3.2 Transcendence with Low-Temperature Sampling
The above shows that, even though transcendence cannot be achieved when directly modeling the distribution, it can be achieved by temperature sampling, assuming that the arg-max predictor achieves higher reward compared to all experts. In other words, we make the subtle connection here that low-temperature sampling can be thought of as performning majority vote [1, 6] between the experts. Please see Appendix A for a formal proof of this connection. When the experts put non-negligible mass onto the best actions, the resulting majority vote may find the best action [9], which improves performance compared to individual experts (i.e., “wisdom of the crowd”) and thus achieve transcendence.
@@ 3.3 Denoising a Single Expert
We now turn to study particular cases where low-temperature sampling can lead to transcendence. The most simple case is of a single expert that outputs a correct but noisy prediction. Denote by $f^{*}$ the optimal expert, s.t. for all $x$ we have  $f^{*}(y|x)=\dfrac{\delta(y\in Y^{*}_{x})}{\lvert Y^{*}_{x}\rvert}$, where $Y^{*}_{x}=\{y\in\mathcal{Y}:y=\max_{y^{\prime}}r(x,y^{\prime})\}$ and $\delta(\text{condition})$ is 1 if the condition is true and 0 otherwise. Now, for some $\rho\in(0,1)$, let $f_{\rho}$ be a “noisy” expert, s.t., for all $x$, with probability $\rho$ chooses a random output, and with probability $1-\rho$ chooses an output according to the optimal expert $f^{*}(\cdot|x)$, namely $f_{\rho}(y|x)=\rho/\left\lvert\mathcal{Y}\right\rvert+(1-\rho)f^{*}(y|x)$. We show that transcendence is achieved with low-temperature sampling for data generated by $f_{\rho}$:
Proposition 3.
@@ 3.4 Transcendence from Multiple Experts
Next, we consider the case where the dataset is generated by multiple experts that complement each other in terms of their ability to correctly predict the best output. For example, consider the case where the input space is partitioned into $k$ disjoint subsets, $\mathcal{X}=\mathcal{X}_{1}\dot{\cup}\dots\dot{\cup}\mathcal{X}_{k}$, s.t. the $i$-th expert performs well on the subset $\mathcal{X}_{i}$, but behaves randomly on other subsets. Namely, assume the expert $f_{i}$ behaves as follows:  $f_{i}(y|x)=\biggl{(}\frac{\delta(y\in Y^{\star}_{x})\delta(x\in\mathcal{X}_{i})}{|Y^{\star}_{x}|}+\frac{\delta(x\notin\mathcal{X}_{i})}{|\mathcal{Y}|}\biggr{)}$ where $Y_{x}^{*}$ is as previously defined and $\delta(\text{condition})$ is 1 if the condition is true and 0 otherwise. We show that, assuming that the test distribution $p_{\mathrm{test}}$ is not concentrated on a single subset $\mathcal{X}_{i}$, we achieve transcendence with low-temperature sampling:
Proposition 4.
@@
In order to build intuition for Proposition 4, see Appendix C for an intuitive diagram.
@@ 4 Experiments
To evaluate the predictive power of our impossibility result of transcendence with no temperature sampling (Proposition 1) as well as our result of transcendence from multiple experts with low temperature sampling (Proposition 2), we turn to modeling and training chess players. Chess stands out as an attractive option for several reasons. Chess is a well-understood domain and more constrained than other settings such as natural language generation, lending to easier and stronger analysis. Evaluation of skill in chess is also natural and well-studied, with several rigorous statistical rating systems available. In this paper, we use the Glicko-2 rating system [7], which is also adopted by https://lichess.org, the free and open-source online chess server from which we source our dataset.
@@ 4.2 Experimental Results
In this section we attempt to answer our primary research question, can low-temperature sampling actually induce transcendence in practice? We test Proposition 2 by evaluating several ChessFormers across different temperature values, from $0.001$ (nearly deterministic), to $1.0$ (original distribution), to $1.5$ (high entropy). In Figure 1 we definitively confirm the existence of transcendence. Our ChessFormer 1000 (where the latter number refers to the maximum rating seen during training) and ChessFormer 1300 models are able to transcend to around 1500 rating at temperature $\tau$ equal to $0.001$. Interestingly, ChessFormer 1500 is unable to transcend at test time, a result we further analyze in Dataset Diversity.
@@
$F(f^{\prime},f;x)=\mathbb{E}_{x\sim d^{f^{\prime}},y\sim f^{\prime}(\cdot|x)}[r(x,y)]-\mathbb{E}_{x\sim d^{f^{\prime}},y\sim f(\cdot|x)}[r(x,y)].$ (4)
@@
In Figure 4, we find that lowering the temperature has the effect of skewing the expected reward distribution to the right, especially for the green $\tau=0.001$ distribution. This result implies that the model does not improve the expected reward by a small amount for many game states, but rather improves the expected reward by a relatively large amount for a few game states. Thus, $\tau=0.001$ improves the expected reward (probability of winning) by an average of $\mathbf{2.15\pm 0.17\%}$, but for some states, this expected improvement is over 5%. Note that the original temperature expected reward can be thought of as a Dirac distribution centered at 00.  The above finding answers our research question in this section: Low-temperature sampling is able improves the expected reward by relatively large amounts for some specific game states, which is likely why the ChessFormer $1000$ and $1300$ model was able to achieve transcendence.
@@
Importantly, we cannot calculate this normalized entropy for every state, as most states after move $16$ in the midgame and before the engame are unique within the dataset and we therefore observe just a single action for thus states. Therefore our metric is limited in that it only considers opening moves, the beginning of the midgame, and the endgame. We consider only common states with greater than $100$ actions by sampling $1,000,000$ games from each dataset. The average entropy confirm our hypothesis: The $<1500$ cut off dataset has on average less diversity than the $<1300$ dataset, which has is again less than the $<1000$ dataset. This result suggests that Chessformer $1500$ likely is not transcendent due to a lack of diversity in its dataset. If the entropy instead stayed constant for each dataset, it would imply that each had a similar level of diversity. In such a case, we would expect that ChessFormer $1500$ likely would also transcend. Instead, as predicted, it is likely not transcendent due to a lack of diversity.
@@
4.3 Additional Settings
SQuADv2 Natural Language Temperature Denoising Experiment. We extend our analysis to the Natural Language Processing domain by running experiments on the Stanford Question Answering Dataset (SQuAD 2.0). We tested the effects of temperature denoising on the performance of several large language models (LLMs) of varying sizes. The SQuAD task involves reading comprehension and question-answering based on Wikipedia articles, making it an ideal setting to evaluate the impact of denoising on language models. We measured the exact-match, semantic-match, and F1 scores of the model outputs at different temperatures. The results show that temperature denoising leads to improved performance, corroborating the findings of our chess experiments and providing broader validation of the underlying mechanism of temperature denoising in diverse domains.
Figure 6: We evaluate several pretrained language models on the SQuADv2 Question-Answering reading comprehesion dataset, a task consisting of answering a question given some snippet from a Wikipedia article. We report F1, ’Exact Match’, and ’Semantic Match’ scores of several different language models of varying size from 163M parameters to 7B parameters, over several different temperatures. Semantic Match is calculated by using another LLM (llama3.1) to judge if two responses are equivalent, even if the exact strings slightly differ between the model output and the correct response. We also report 95% confidence intervals calculated through taking ±$\pm 1.96\sigma$.
Toy Model Setting and Results. In addition, we develop a toy theoretical model to further study when transcendence is possible. This model involves a classification task with Gaussian input data and linearly separable classes. Experts label the data with noisy versions of the ground truth separator. We trained a linear model on a dataset labeled by random experts and observed the test accuracy for different temperature settings. The synthetic experiments demonstrated that transcendence occurs when expert diversity is high and temperature is low, aligning with our theoretical and empirical analysis in the chess domain.
Figure 7: Toy model for demonstrating transcendence. Input data is $d$-dimensional Gaussian, with $d=100$. Output is classification with $10$ classes. Ground-truth is generated by a linear function, i.e. $y=\arg\max_{i}W_{i}^{\star}x$ for some $W^{*}\in\mathbb{R}^{10\times d}$. We sample $k$ experts, with $k=5$, to label the data, where the labels of each expert are generated by some $W\in\mathbb{R}^{10\times d}$ s.t. $W=W^{*}+\xi$, where $\xi_{i,j}\sim\mathcal{N}(0,\sigma^{2})$, for some standard deviation $\sigma$. Namely, each expert labels the data with a noisy version of the ground truth separator, with noise std $\sigma$. We then train a linear model on a dataset with $10K$ examples, where each example is labeled by a random expert. We plot the test accuracy, measured by the probability assigned to the correct class, for different choices of temperature, and compare to the best expert.
@@ 5 Related Work
Another historical thread in AI research is the strength of diverse learners. Long since the development of ensemble methods that exploit learner diversity—including bagging [1], boosting [6], and model averaging [19]—researchers have continued to articulate this insight across settings. Similar to our chess setting, a diverse team of go playing agents have been proven and empirically shown to outperform solitary agents [9] and homogeneous teams [28], even when the alternative models individually outperform the diverse team members [17]. We draw a connection to this deep literature through our theory, which shows that imitation learning objective and then performing low-temperature sampling subtly implies the same principle of majority voting. Teacher diversity has also been explored in the machine learning literature. One related method is ensemble distillation [16], in which a model is trained with an additional objective to match a variety of weaker teacher models. Closer to our setting, ensemble self-training approaches [24] train a learner directly on the labels produced by varied teachers. Large language models supervised by smaller or less trained models are said to exhibit “weak to strong generalization” [2]. Overall, evidence continues to accrue that the general phenomenon we address is pervasive: that is, models can substantially improve over the experts that generate their training data.
@@
Our work also draws connections to the Offline Reinforcement Learning [14] setting, where one attempts to learn a new policy $\pi$ that improves upon a fixed dataset generated by some behavior policy $\pi_{\beta}$. However, our setting of imitation learning differs substantially from this literature, as we do not explicitly train our model on a RL objective that attempts to improve upon the dataset. Importantly, such an objective oftentimes introduces training instabilities [15] and also assumes reward labels. We defer a more extended discussion of related work to Appendix D.
@@ 6 Discussion and Future Work
This paper introduces the concept of transcendence. Our theoretical analysis shows that low-temperature sampling is key to achieving transcendence by denoising expert biases and consolidating diverse knowledge. We validate our findings empirically by training several chess models which, under low-temperature sampling, surpass the performance of the players who produced their training data, as well as further experiments in natural language question-answering and toy Gaussian models. We additionally highlight the necessity of dataset diversity for transcendence, emphasizing the role of varied expert perspectives.
@@
The possibility of “superintelligent” AGI has recently fueled many speculative hopes and fears. It is therefore possible that our work will be cited by concerned communities as evidence of a threat, but we would highlight that the denoising effect addressed in this paper does not offer any evidence for a model being able to produce novel solutions that a human expert would be incapable of devising. In particular, we do not present evidence that low temperature sampling leads to novel abstract reasoning, but just denoising of errors.
@@ Appendix A Proofs
Here we prove Proposition 1, where transcendencecannot occur by purely using imitation learning in our setting where all experts are sampled uniformly across the input distribution.
@@
We now give the proof of Proposition 2 that if the arg-max prediction is better than the best expert, then transcendence is possible with low-temperature sampling.
@@
To prove Proposition 3, we directly use the result in Proposition 2.
@@
Next, we give the proof for that low-temperature sampling can be thought of as performing majority vote [1, 6] between the experts:
Proposition 5.
Let $\mathbf{z}=[z_{1},z_{2},\dots,z_{n}]$ be a vector and $\tau>0$ be a temperature parameter. Define the softmax function as
$\sigma_{\tau}(z_{i})=\frac{e^{z_{i}/\tau}}{\sum_{j=1}^{n}e^{z_{j}/\tau}}.$
Then, as $\tau\to 0^{+}$, the limit of the softmax function is given by
$\lim_{\tau\to 0^{+}}\sigma_{\tau}(z_{i})=\begin{cases}\displaystyle\frac{1}{k},&\text{if }z_{i}=z_{\max},\\ 0,&\text{otherwise},\end{cases}$
where $z_{\max}=\max_{1\leq j\leq n}z_{j}$, and $k$ is the number of indices $i$ such that $z_{i}=z_{\max}$.
Proof.
Let $z_{\max}=\max_{1\leq j\leq n}z_{j}$, and define the set
$S=\{i\mid z_{i}=z_{\max}\},$
with cardinality $k=|S|$.
For each $i$, let
$\Delta_{i}=z_{i}-z_{\max}\leq 0.$
Then the softmax function becomes
$\sigma_{\tau}(z_{i})=\frac{e^{(z_{\max}+\Delta_{i})/\tau}}{\sum_{j=1}^{n}e^{(z_{\max}+\Delta_{j})/\tau}}=\frac{e^{\Delta_{i}/\tau}}{\sum_{j=1}^{n}e^{\Delta_{j}/\tau}},$
since $e^{z_{\max}/\tau}$ cancels out in the numerator and denominator.
We analyze the behavior of the terms as $\tau\to 0^{+}$.
For $i\in S$ we have $\Delta_{i}=0$ and so:
$e^{\Delta_{i}/\tau}=e^{0}=1.$
For $i\notin S$ we have $\Delta_{i}<0$ so
$\lim_{\tau\to 0}e^{\Delta_{i}/\tau}=0$
Therefore, the denominator simplifies to
$\lim_{\tau\to 0^{+}}\sum_{j=1}^{n}e^{\Delta_{j}/\tau}=\sum_{j\in S}\lim_{\tau\to 0^{+}}e^{\Delta_{j}/\tau}+\sum_{j\notin S}\lim_{\tau\to 0^{+}}e^{\Delta_{j}/\tau}=\sum_{j\in S}1+\sum_{j\notin S}0=k.$
Similarly, the numerator becomes
$\lim_{\tau\to 0^{+}}e^{\Delta_{i}/\tau}=\begin{cases}1,&\text{if }i\in S,\\ 0,&\text{if }i\notin S.\end{cases}$
Thus, for each $i$,
$\lim_{\tau\to 0^{+}}\sigma_{\tau}(z_{i})=\frac{\lim_{\tau\to 0^{+}}e^{\Delta_{i}/\tau}}{\lim_{\tau\to 0^{+}}\sum_{j=1}^{n}e^{\Delta_{j}/\tau}}=\begin{cases}\displaystyle\frac{1}{k},&\text{if }i\in S,\\ 0,&\text{if }i\notin S.\end{cases}$
This concludes the proof. ∎
Finally, we give the proof of Proposition 4, or the statement that transcendence can occur from multiple experts if the test distribution $p_{\mathrm{test}}$ is spread across multiple disjoing subsets of $\mathcal{X}_{i}$.
@@
Thus, we get $f_{\max}=f^{*}$, and the required follows from Proposition 2.